OpenStudy (anonymous):

let f(x) = x + 4 / 5. Find f^-1 (6)

OpenStudy (anonymous):

That notation up there means we need to get the inverse of the function. In doing this, we first isolate x. And make the xs y and the ys x. y=x+4 y-4=x x-4=y f-1(x) = x-4 f^-1(6) = 6-4=2.